A boy in a stationary boat on a still pond has lost his oars in water. In order to get the boat moving again, he throwa his rucksack horizontally out of boat with speed of 4m/s. Mass of boat is 60kg, mass of boy is 40kg mass of rucksack is 5kg. How fast will the action make the boat move?
1
Expert's answer
2021-10-27T08:48:04-0400

The key for solution is the conservation of momentum: the initial momentum is zero, thus, the final momentum must be zero as well:


"0=m_rv_r-(m_b+M_b)v_b,\\\space\\
v_b=\frac{m_rv_r}{m_b+M_b}=\frac{5·4}{40+60}=0.2\text{ m/s}."